Find the strain of the 2.0 mm copper wire that is subject to a 1.0kN tensile force.
1
Expert's answer
2020-11-30T14:53:47-0500
"\epsilon=\frac{\sigma}{Y}\\\epsilon=\frac{F}{\pi r^2Y}\\\epsilon=\frac{1000}{\pi (0.002)^2(117\cdot10^9)}=0.00068"